OpenStudy (anonymous):

use the logarithmic differentiation: y = x-sin(x)....i got it to be.... x^(sinx)(sinx/x+lnxcosx)....im not sure if thats right...

OpenStudy (anonymous):

Why would you ever be asked to do that with logarithmic?

OpenStudy (anonymous):

thats what the assignments ask

OpenStudy (anonymous):

which i have to derviative

OpenStudy (anonymous):

Logarithmic just makes that complicated. Its supposed to simplify complicated stuff. You could do: ln(y)=ln(x-sin(x)) y'/y=(1-cos(x))/(x-sin(x)) y'=y((1-cos(x))/(x-sin(x)) Replacing y. You are left with just y'=1-cos(x)

OpenStudy (anonymous):

is it \[x^{-\sin(x)}\]?

OpenStudy (anonymous):

its \[y=x ^{sinx}\]

OpenStudy (anonymous):

i got it to be \[y=x ^{sinx}(sinx/x+lnxcosx)....\]

OpenStudy (anonymous):

AHHHHH lol. ln(y)=sin(x)ln(x) y'/y=cos(x)ln(x)+sin(x)/x y'=y(cos(x)ln(x)+sin(x)/x) y'=x^(sin(x))(cos(x)ln(x)+sin(x)/x)

OpenStudy (anonymous):

but where did i go wrong?

OpenStudy (anonymous):

You have the same thing I do. You did it fine, you just do your product rule in a different order :P

OpenStudy (anonymous):

o ok...well i guess im getting better

OpenStudy (anonymous):

good work!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

thanks guys : )